Circular No. 9010

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) CBAT@CFA.HARVARD.EDU (science)

        URL http://www.cfa.harvard.edu/iau/cbat.html ISSN 0081-0304

Phone 617-495-7440/7244/7444 (for emergency use only)

COMETS 210P/CHRISTENSEN AND 211P/HILL Comet P/2008 X4 = P/2003 K2 (Christensen; cf. IAUCs 9005, 9007,

9008) has been assigned the permanent designation 210P. Comet P/2008 X1 = P/2003 F6 (Hill; cf. IAUCs 9001, 9009) has been assigned the permanent designation 211P. See also MPC 64693.

COMET 212P/2000 YN_{30} (NEAT) An apparent asteroidal object discovered by the NEAT survey (the dis-

covery observation tabulated below was originally assigned the designation 2000 YN_{30} and published on MPO 24160) was recovered a few months ago with the 0.9-m Spacewatch reflector at Kitt Peak (recovery observation also tabulated below; cf. MPEC 2008-V41 and MPO 263011) but only recently noted to show cometary activity. A. R. Gibbs reports that four co-added 60-s unfiltered CCD exposures taken on 2009 Jan. 2.53 with the Catalina 0.68-m Schmidt telescope show a compact coma 7′′ across with a narrow 35′′ tail in p.a. 285^{o}. Y. C. Cheng and H. Y. Hsiao, Institute of Astronomy, National Central University, Jhongli, Taiwan, report that three 300-s Rband images obtained on Jan. 2.8 with the Lulin 1-m telescope show total mag 18.2–18.4 and a tail < 20′′ long in p.a. 290^{o}. J. V. Scotti writes that fifteen co-added exposures with the Spacewatch 1.8-m reflector on Jan. 7.4 show an 8′′ coma and a narrow tail extending at least 2′.2 in p.a. 283^{o}; the tail’s width expands with distance from the nuclear condensation from ∼ 8′′ to 12′′ as it fades.

Date UT α_{2000} δ_{2000} Mag. Observer
2000 Dec. 26.46926 4^{h}19^{m}20.^{s}45 + 6^{o}03′33′′.1 20.4 NEAT
2008 Oct. 30.48499 8 50 58.52 +15 37 57.6 20.5 Spacewatch

The following orbital elements by B. G. Marsden, Smithsonian Astrophysical Observatory, from 69 observations, 2000-2009 (mean residual 0′′.6), appear on MPC 64768 (with new observations on MPC 64748-64749). The permanent designation 212P has been assigned to this comet.

Epoch = 2008 Nov. 30.0 TT

T = 2008 Dec. 3.2698 TT  ω = 15.^{o}0488
e = 0.578876  Ω = 98.9290 }2000.0
q = 1.654465 AU  i = 22.3979

a = 3.928686 AU n^{o} = 0.1265707 P = 7.79 years

2009 January 12  \begin{matrix}c \\ ◯ \end{matrix} Copyright 2009 CBAT  Daniel W. E. Green